OpenStudy (anonymous):

i have a total of 49 dimes and quarters. the value of quarters is the same value of my dimes. find the number of each coin

OpenStudy (anonymous):

what are my equations?

OpenStudy (whpalmer4):

okay, there are 49 coins, let's call the variables representing quarters and dimes \(q\) and \(d\) respectively. \[q+d = 49\] Agreed?

OpenStudy (anonymous):

I got that part (:

OpenStudy (anonymous):

Just the second equation that is stumping me

OpenStudy (whpalmer4):

no problem. what's the value of \(q\) quarters (expressed in cents)?

OpenStudy (whpalmer4):

1 quarter = 25 cents 2 quarters = 50 cents \(q\) quarters = ?

OpenStudy (anonymous):

25q...?

OpenStudy (whpalmer4):

yes. what is the corresponding expression for the value of the dimes?

OpenStudy (anonymous):

10d

OpenStudy (whpalmer4):

right. now, the problem says the value of the quarters is the same as the value of the dimes, right? How could you express that?

OpenStudy (anonymous):

25q=10d

OpenStudy (whpalmer4):

okay, so now you have two equations in two unknowns, and as one of them has coefficients of 1 for both variables, solving by substitution should be easy enough...

OpenStudy (anonymous):

Ohhhhh ok. THANKS SO MUCH! (:

OpenStudy (whpalmer4):

got an answer for me?

OpenStudy (anonymous):

oh ok hold on

OpenStudy (anonymous):

Wait....wouldn't i have to divide the second equation by 25 or 10 for the variable to be on it's own?

OpenStudy (anonymous):

14 quarters, 35 dimes

OpenStudy (whpalmer4):

14*25 + 35*10 = 350 + 350 = 700 so that is correct equal value of quarters and dimes, and 14+35 = 49 coins in all.